Write a system of two equations in two unknowns. Solve the system by using the substitution method. « Bill and Sue together earn $76,433 per year. If Bill earns $15,000 less than twice Sue’s salary, then how much does each person earn per year? »

Answer by checkley71(8403)

76433=(2x-15000)+x or 3x=91433 or x=91433/3 or x=30477.67 is Sue's salary &
2*30477.67-15000=45955.33 proof 30477.67+45955.33=76433.00